OpenStudy (jojo12g):

i think a What is the value of the expression k – 14 when k = –17? A. –31 B. –3 C. 3 D. 31

OpenStudy (thread252):

and you are correct! it is A

OpenStudy (jojo12g):

thanks

OpenStudy (lochana):

@thread252 correct:)

OpenStudy (trojanpoem):

Just plug in the value given for the variable k , k = -17 let f(k) = k -14 , when k = -17 f(-17) = -17- 14 = -31(A) as thread252 said.